# A Low Rank Neural Representation of Entropy Solutions

### v1

@@ 5.2 A time-layered approximation of the rarefied characteristics
We first introduce an approximation of $\widehat{X}(x,t)$ Eq. 2.22 that is accurate for a given time interval. This approximation will be uniformly accurate, with the error controlled by the number of degrees of freedom depending on the complexity of $u_{0}$ and the flux function $F$.
@@
Note that these endpoints can represent arbitrary number of shocks; the ranges ${\vdash}(\Upsilon_{t},t)$ and ${\dashv}(\Upsilon_{t},t)$ yield left and right endpoints of all shocks that have formed up to time $t$. For the points $x\in\Upsilon_{t}$, that is, the points in $\widehat{\Omega}_{x}$ whose subsequent characteristic curves go into the shock at time $t$, the function values ${\vdash}(x,t)$ and ${\dashv}(x,t)$ are the left and right most end points of the maximal interval containing the point that are in the shock at the point.
@@
Furthermore, since we assumed our initial condition was in $\overline{\mathcal{U}}$ Eq. 2.3, and our problem is homogeneous Eq. 2.1,
@@
By construction we have for any $t_{k}\in\Omega_{t}$
(5.14) $\widehat{X}(\cdot,t_{k})=\Xi_{k}(\cdot,t_{k}),$
and for $t\in[t_{k},t_{k+1}]$
(5.15) $\widehat{X}(z,t)=\Xi_{k}(z,t)\quad\text{ for all }z\in\widehat{\Omega}_{x}\setminus\Upsilon_{t}.$
Let us fix $x_{0}\in\Upsilon_{t_{k}}$ and recall $\widehat{X}_{0}$ is the rarefied characteristics with multivalued inverse Eq. 2.27. Then we have
(5.16) $E_{k}(t;x_{0})$ $:=\sup_{x\in[{\vdash}(x_{0},t),{\dashv}(x_{0},t)]}\lVert{\widehat{X}(x,t)-\Xi_{k}(x,t)}\rVert$
where we let
(5.17) $\overline{u}_{\text{loc}}$ $:=\sup_{\begin{subarray}{c}z\in[{\vdash}(x_{0},t),{\vdash}(x_{0},t_{k})]\\ \qquad\cup[{\dashv}(x_{0},t_{k}),{\dashv}(x_{0},t)]\end{subarray}}u_{0}(z),\quad$
Since $[\overline{u}_{\text{loc}},\underline{u}_{\text{loc}}]\subset u_{0}(\Omega_{x})$, we obtain the first estimate Eq. 5.12.
(5.18) $\left\lvert\Xi_{k}(\cdot,t)\right\rvert_{TV(\widehat{\Omega}_{x})}$ $\leq\left\lvert\Xi_{k}(\cdot,t)\right\rvert_{TV(\Upsilon_{t_{k}})}+\left\lvert\Xi_{k}(\cdot,t)\right\rvert_{TV(\widehat{\Omega}_{x}\setminus\Upsilon_{t_{k}})}$
(5.19) $\left\lvert\Xi_{k}(\cdot,t)\right\rvert_{TV([{\vdash}(x_{0},t_{k}),{\dashv}(x_{0},t_{k})])}$ $=\left\lvert\widehat{X}_{0}({\vdash}(x_{0},t_{k}),t)-\widehat{X}_{0}({\dashv}(x_{0},t_{k}),t)\right\rvert$
(5.20) $\lvert{\Xi_{k}(\cdot,t)}\rvert_{TV(\Upsilon_{t_{k}})}\leq\lvert{\widehat{X}_{0}(\cdot,t)}\rvert_{TV(\Upsilon_{t_{k}})}.$
Using this in Eq. 5.18, we have that
(5.21) $\left\lvert\Xi_{k}(\cdot,t)\right\rvert_{TV(\widehat{\Omega}_{x})}$ $\leq\lvert{\widehat{X}_{0}(\cdot,t)}\rvert_{TV(\widehat{\Omega}_{x})}=\lvert{\hat{I}+tF^{\prime}\circ\hat{u}_{0}}\rvert_{TV(\widehat{\Omega}_{x})}$
Figure 7: An illustration of time-layered approximation $\Xi_{k}$ Eq. 5.10.
Figure 8: An illustration of the interpolant $\eta_{k}$ Eq. 5.28.
Note that the upper bound Eq. 5.12 with the TV norm $\left\lvert{u_{0}}\right\rvert_{TV(\Omega_{x})}$ obtained from the last term in Eq. 5.16 is not sharp, and can be replaced by $2\lVert{u_{0}}\rVert_{L^{\infty}(\Omega_{x})}$ because
@@ 5.3 The spatial discretization of time-layered approximation
Now we introduce the appropriate spatial discretization of the time-layered approximations $\Xi_{k}$ Eq. 5.10 that we have constructed in the previous section.
@@
In the following, we will make use of the estimates Eqs. 5.12 and 5.13 in Lemma 5.1, and the uniform time grid $t_{k}$ with grid size $K$ Eq. 5.23.
Because of the TV bound Eq. 5.13, there is an adaptive grid $(z_{i})_{i=1}^{K}$ over $\widehat{\Omega}_{x}$ on which we have an approximation $\barwedge_{k}(\cdot,t)\in C(\widehat{\Omega}_{x}\times\Omega_{t})\cap P_{1}(\widehat{\Omega}_{x}\times\Omega_{t})$ of $\Xi_{k}(\cdot,t)$ satisfying
@@
(5.27) $\barwedge_{k}(z_{i})=\left\{\begin{aligned} \widehat{X}_{0}(z_{i},t_{k})&\quad\text{ for }z_{i}\in\widehat{\Omega}_{x}\setminus\Upsilon_{t},\\ \widehat{X}(z_{i},t_{k})&\quad\text{ for }z_{i}\in\Upsilon_{t}.\end{aligned}\right.$
We have now obtained the interpolants $(\barwedge_{k})$, and next we wish to exploit the fact that these interpolants are closely related to each other across the index $k$. They interpolate $(\Xi_{k})$ each in the form Eq. 5.10 containing the functions $(v_{k})$. Observe that these functions approximate $F^{\prime}\circ\hat{u}_{0}$ in $\widehat{\Omega}_{x}\smallsetminus\Upsilon_{t_{k}}$. Since $(v_{k})$ each approximate the same spatial function, it is possible to construct an approximation of $(\Xi_{k})$ for all $k$ simply by reorganizing the point values of $(\barwedge_{k})$ by adding or removing the differences $(v_{k+1}-v_{k})$. We do this by introducing a continuous piecewise linear interpolant of these differences.
@@
due to Eq. 5.12 and Eq. 5.26.
Next, let us set the function $\barwedge:\widehat{\Omega}_{x}\times\Omega_{t}\to\mathbb{R}$ which is continuous piecewise linear in space but only piecewise linear in time,
(5.30) $\barwedge(x,t):=\barwedge_{K}(x)+t\left[\sum_{k=0}^{K-1}\eta_{k}(x)\right]$
@@
(5.32) $\doublebarwedge(x,t):=\barwedge_{K}(x,t)+t\left[\sum_{k=0}^{K-1}\eta_{k}\left(x-C_{x}\sigma\left(\frac{K}{T}(t-t_{k})\right)\right)\right],$
for a constant $C_{x}\geq\left\lvert\Omega_{x}\right\rvert$.
@@
(5.33) $\doublebarwedge(x,t)\approx\barwedge_{K}(x,t)+t\left[\sum_{k=0}^{K-1}\eta_{k}\left(x-C_{x}\sigma\left(\frac{K}{T}(t-\lambda(z_{i_{k}}))\right)\right)\right],$
@@
The approximation $\doublebarwedge\in C(\widehat{\Omega}_{x}\times\Omega_{t})\cap P_{1}(\widehat{\Omega}_{x}\times\Omega_{t})$ defined above in Eq. 5.32 satisfies
@@
By construction, $\doublebarwedge(\cdot,t)$ is based on $\eta_{k}\in C(\widehat{\Omega}_{x})\cap P_{1}(\widehat{\Omega}_{x})$ given in Eq. 5.28 for index $k=1,...\,,K-1$ in Eq. 5.32 so we have for $t\in[t_{k},t_{k+1}]$
(5.35) $\left\lVert{\widehat{X}(\cdot,t)-\doublebarwedge(\cdot,t)}\right\rVert_{L^{\infty}(\widehat{\Omega}_{x})}$
upon applying Eq. 5.29 and Eq. 5.31. The RHS does not depend on the interval $[t_{k},t_{k+1}]$ that $t$ belongs to, so the result follows.
Figure 9: A diagram depicting the approximation by $\doublebarwedge(x,t)$.
@@
Let $u$ be the entropy solution to the scalar conservation law Eq. 2.1. It is known (see, e.g. [18]) that there is an entropy solution $\bar{u}$ with piecewise constant initial datum $\bar{u}_{0}\in\overline{\mathcal{U}}$ satisfying
Let $\hat{u}_{0}$ be the extended initial condition of such $\bar{u}_{0}$ constructed as in Eq. 2.20. Let $\hat{u}_{0\varepsilon}$ be its approximation of the form
then it is known (see, e.g. [22]) that such a $\hat{u}_{0\varepsilon}$ can be chosen to satisfy
[…]
@@
(5.41) $\widehat{\Omega}_{x}=(\widehat{\Omega}_{x}\setminus\Upsilon_{t})\cup\Upsilon_{t},$
we have that $\widehat{X}$ is constant in $\Upsilon_{t}$ so that $\lVert{\widehat{X}^{\prime}}\rVert_{L^{\infty}(\Upsilon_{t})}=0$. Now recall the intervals inserted into in the extended domain $\Gamma_{x}=\bigcup_{k=1}^{n_{\mathcal{R}}}\Gamma_{k}$ defined in Eq. 2.18. Writing the shock-free part of the extended domain $\widehat{\Omega}_{x}\setminus\Upsilon_{t}$ into $\Gamma_{x}$ and its complement
(5.42) $\widehat{\Omega}_{x}\setminus\Upsilon_{t}=\left((\widehat{\Omega}_{x}\setminus\Upsilon_{t})\,\setminus\Gamma_{x}\right)\cup\Gamma_{x},$
we straightforwardly bound,
(5.43) $\lVert{\widehat{X}^{\prime}}\rVert_{L^{\infty}(\widehat{\Omega}_{x})}$ $\leq\lVert{\widehat{X}^{\prime}}\rVert_{L^{\infty}(\widehat{\Omega}_{x}\setminus\Upsilon_{t})}+\lVert{\widehat{X}^{\prime}}\rVert_{L^{\infty}(\Upsilon_{t})}$
Due to the way of $\hat{u}_{0}$ was defined in Eq. 2.20, for $k=1,...\,,n_{\mathcal{R}}$,
(5.44) $\lVert{(F^{\prime}\circ\hat{u}_{0})^{\prime}}\rVert_{L^{\infty}(\Gamma_{k})}$ $=\left\lVert{\nu_{k}^{-}+\frac{\nu_{k}^{+}-\nu_{k}^{-}}{\varepsilon_{k}}(\cdot-\iota(\xi_{k}))}\right\rVert_{L^{\infty}(\iota(\xi_{k})+[0,\varepsilon_{k}))}$
(5.45) $\lVert{\widehat{X}^{\prime}}\rVert_{L^{\infty}(\widehat{\Omega}_{x})}\leq 1+T\lVert{(F^{\prime}\circ u_{0})^{\prime}}\rVert_{L^{\infty}(\Omega_{x})}\leq 1+T\lVert{F^{\prime\prime}}\rVert_{L^{\infty}(u_{0}(\Omega_{x}))}\left\lvert{u_{0}}\right\rvert_{TV(\Omega_{x})},$
which is a generalization of the crude bound Eq. 4.43 for classical characteristics.
Due to Lemma 4.3, we have
(5.46) $\lVert{u(\cdot,t)-\bar{h}(\cdot,t)}\rVert_{L^{1}(\Omega_{x})}$
The result follows from the estimates Eq. 5.34 and Eq. 5.45, and further,
(5.47) $\lVert{u(\cdot,t)-\bar{h}(\cdot,t))}\rVert_{L^{1}(\Omega_{x})}$
Gathering the terms,
(5.48) $\lVert{u(\cdot,t)-\bar{h}(\cdot,t))}\rVert_{L^{1}(\Omega_{x})}$
The approximation error in Eq. 5.40 scales as $\sim K^{-1}$ and the constants depend only on the initial condition and the flux function. is important to note that this rate implies that our construction is not a naïve space-time approximation, for which the error would scale with the d.o.f $\sim K^{-\frac{1}{2}}$.
@@
Moreover, in this space-time approximation, the temporal approximation is trivial, due to the explicit linear time dependence as clearly seen in Eq. 5.32.
@@ 5.4 The LRNR approximation of compositional form
There is a LRNR $H_{r}$ with rank $r=2$, d.o.f $K$, width $\sim K^{2}$, and depth $L=5$, such that for the solution to Eq. 2.1 with initial condition $u_{0}\in\mathcal{U}$ at each time $u(t):=u(\cdot,t)$ there is a $h(t):=h(\gamma(t),\theta(t))\in H_{r}$ with the approximation error
@@
where the coefficients $\gamma(t),\theta(t)$ are affine functions of $t$.
Note that neither the rank $r$ nor the number of layers $L$ depends on the initial condition. This result shows there is an upper bound on a kind of “nonlinear width” as described by the number of coefficients $r$ and the number of layers $L$, namely the tuple $(r,L)=(2,5)$ associated with the entropy solutions of the scalar conservation law Eq. 2.1 regardless of the complexity of the initial data or the shock topology structure.

### v2

@@ 5.2 A time-layered approximation of the rarefied characteristics
We first introduce an approximation of $\widehat{X}(x,t)$ (2.22) that is accurate for a given time interval. This approximation will be uniformly accurate, with the error controlled by the number of degrees of freedom depending on the complexity of $u_{0}$ and the flux function $F$.
@@
Note that these endpoints can represent arbitrary number of shocks; the ranges ${\vdash}(\Upsilon_{t},t)$ and ${\dashv}(\Upsilon_{t},t)$ yield left and right endpoints of all shocks that have formed up to time $t$. For the points $x\in\Upsilon_{t}$, that is, the points in $\widehat{\Omega}_{x}$ whose subsequent characteristic curves go into the shock at time $t$, the function values ${\vdash}(x,t)$ and ${\dashv}(x,t)$ are the left and right most end points of the maximal interval containing the point that are in the shock at the time.
@@
Furthermore, since we assumed our initial condition was in $\overline{\mathcal{U}}$ (2.3), and our problem is homogeneous (2.1),
@@
See Appendix B.4.
Figure 8: An illustration of time-layered approximation $\Xi_{k}$ (5.12) immediately after shock formation, compared with the rarefied characteristics with multivalued inverse $\widehat{X}_{0}$ and the rarefied characteristics $\widehat{X}$. At uniform grid times $t_{k}$ (5.17) the $\Xi_{k}(\cdot,t_{k})=\widehat{X}(\cdot,t_{k})$, however, as the time is evolved forward $\Xi_{k}(\cdot,t)$ is different from $\widehat{X}(\cdot,t_{k})$ in general, and the difference is within the $\epsilon$ error bound.
Figure 9: An illustration of the interpolant $\eta_{k}$ (5.22) along with the functions $v_{k}$ and $v_{k+1}$ (5.13). The thin slice $\eta_{k}$ is plotted to the right.
Note that the upper bound (5.14) with the TV norm $\left\lvert{u_{0}}\right\rvert_{TV(\Omega_{x})}$ obtained from the last term in (B.25) is not sharp, and can be replaced by $2\lVert{u_{0}}\rVert_{L^{\infty}(\Omega_{x})}$ because
@@ 5.3 The spatial discretization of time-layered approximation
Now we introduce the appropriate spatial discretization of the time-layered approximations $\Xi_{k}$ (5.12) that we have constructed in the previous section.
@@
In the following, we will make use of the estimates (LABEL:eq:hXZk_error,eq:ZkTV) in Lemma 5.1, and the uniform time grid $t_{k}$ with grid size $K$ (5.17).
Because of the TV bound (5.15), there is an adaptive grid $(z_{i})_{i=1}^{K}$ over $\widehat{\Omega}_{x}$ on which we have an approximation $\barwedge_{k}(\cdot,t)\in C(\widehat{\Omega}_{x}\times\Omega_{t})\cap P_{1}(\widehat{\Omega}_{x}\times\Omega_{t})$ of $\Xi_{k}(\cdot,t)$ satisfying
@@
(5.21) $\barwedge_{k}(z_{i})=\left\{\begin{aligned} \widehat{X}_{0}(z_{i},t_{k})&\quad\text{ for }z_{i}\in\widehat{\Omega}_{x}\setminus\Upsilon_{t_{k}},\\ \widehat{X}(z_{i},t_{k})&\quad\text{ for }z_{i}\in\Upsilon_{t_{k}}.\end{aligned}\right.$
We have now obtained the interpolants $(\barwedge_{k})$, and next we wish to exploit the fact that these interpolants are closely related to each other across the index $k$. They interpolate $(\Xi_{k})$ each in the form (5.12) containing the functions $(v_{k})$. Observe that these functions approximate $F^{\prime}\circ\hat{u}_{0}$ in $\widehat{\Omega}_{x}\smallsetminus\Upsilon_{t_{k}}$. Since $(v_{k})$ each approximate the same spatial function, it is possible to construct an approximation of $(\Xi_{k})$ for all $k$ simply by reorganizing the point values of $(\barwedge_{k})$ by adding or removing the differences $(v_{k+1}-v_{k})$. We do this by introducing a continuous piecewise linear interpolant of these differences.
@@
due to (5.14) and (5.20).
Next, let us set the function $\barwedge:\widehat{\Omega}_{x}\times\Omega_{t}\to\mathbb{R}$ which is continuous piecewise linear in space but only piecewise linear in time. Let $\barwedge_{K}(x,t):=\hat{I}(x)+\frac{t}{T}(\barwedge_{K}(x)-\hat{I}(x))$
(5.24) $\barwedge(x,t):=\barwedge_{K}(x,t)+t\left[\sum_{k:t<t_{k}}\eta_{k}(x)\right]$
@@
(5.26) $\doublebarwedge(x,t):=\barwedge_{K}(x,t)+t\sum_{k=0}^{K-1}\eta_{k}\left(x-C_{x}\sigma\left(\frac{K}{T}(t-t_{k})\right)\right),$
for a constant $C_{x}\geq\left\lvert\Omega_{x}\right\rvert$. Notice how the summation indices are over all $k$: Essentially, through the modification $x\mapsto x-C_{x}\sigma\left(\frac{K}{T}(t-t_{k})\right)$ we are emulating the partial sum over $k$ in (5.24) by quickly translating the $\eta_{k}$ terms outside of the spatial domain $\widehat{\Omega}_{x}$.
@@
(5.27) $\doublebarwedge(x,t)\approx\barwedge_{K}(x,t)+t\sum_{k=0}^{K-1}\eta_{k}\left(x-C_{x}\sigma\left(\frac{K}{T}(t-\lambda(z_{i_{k}}))\right)\right),$
@@
The approximation $\doublebarwedge\in C(\widehat{\Omega}_{x}\times\Omega_{t})\cap P_{1}(\widehat{\Omega}_{x}\times\Omega_{t})$ defined above in (5.26) satisfies
@@
See Appendix B.6.
[…]
Let $\hat{u}_{0}$ be the extended initial condition of such $\bar{u}_{0}$ constructed as in (2.20). Let $\hat{u}_{0\varepsilon}$ be its approximation (see, e.g. [23]) of the form
that satisfies the error bound
@@
See Appendix B.7.
@@
The approximation error in (5.33) scales as $\sim K^{-1}$ and the constants depend only on the initial condition and the flux function. is important to note that this rate implies that our construction is not a naïve space-time approximation, for which the error would scale with the d.o.f $\sim K^{-\frac{1}{2}}$.
[…]
Moreover, in this space-time approximation, the temporal approximation is trivial, due to the explicit linear time dependence as clearly seen in (5.26).
@@ 5.4 The LRNR approximation of compositional form
There is a LRNR $H_{r}$ with rank $r=2$, d.o.f $K$, width $\sim K^{2}$, and depth $L=5$, such that for the solution to (2.1) with initial condition $u_{0}\in\mathcal{U}$ at each time $u(t):=u(\cdot,t)$ there is a $h(t):=h(\gamma(t),\theta(t))\in H_{r}$ with the approximation error
@@
where the coefficients $\gamma(t),\theta(t)$ are linear or constant functions of $t$.
@@
See Appendix B.8.
Note that neither the rank $r$ nor the number of layers $L$ depends on the initial condition. This result shows there is an upper bound on a kind of “nonlinear width” as described by the number of coefficients $r$ and the number of layers $L$, namely the tuple $(r,L)=(2,5)$ associated with the entropy solutions of the scalar conservation law (2.1) regardless of the complexity of the initial data or the shock topology structure.
Acknowledgements
Appendix A Glossary of definitions and relations
Formulations of the entropy solution (Section 2).
  1. A-1.
    $u$ entropy solution (Sec. 2.1)
  2. A-2.
    $u_{0}$ initial condition to the entropy solution (Sec. 2.1)
  3. A-3.
    $\xi$ mapping for similarity solution (2.9)
  4. A-4.
    $X$ characteristic curves (2.5)
    (2.11) $u(x,t)=\left\{\begin{aligned} u_{0}(X^{+}(x,t))\quad&(x,t)\in\Omega\smallsetminus\mathcal{R}\\ (F^{\prime})^{-1}\left(\xi(x,t)\right)\quad&(x,t)\in\mathcal{R}\end{aligned}\right.$
  5. A-5.
    $\hat{u}_{0}$ an extension of the initial condition $u_{0}$ (2.20)
  6. A-6.
    $\widehat{X}$ rarefied characteristic curves (2.22)
    (2.25) $u(x,t)=\hat{u}_{0}\circ\widehat{X}^{+}(x,t)\quad(x,t)\in\Omega_{x}$
  7. A-7.
    $\widehat{X}_{0}$ rarefied chars. with multivalued inverse
    (2.26) $\widehat{X}_{0}(x,t):=\hat{I}(x)+t(F^{\prime}\circ\hat{u}_{0})(x)$
  8. A-8.
    $\rho$ threshold function (2.31)
  9. A-9.
    $\lambda$ shock time function (Sec. 2.3)
  10. A-10.
    $\widecheck{X}$ relief characteristic curves (2.32)
    (2.33) $u(x,t)=\hat{u}_{0}\circ\widecheck{X}^{+}(x,t)\quad(x,t)\in\Omega$
Low rank neural representation (LRNR) (Section 3).
  1. A-11.
    $h(x;W,B)$ feedforward neural network (Def. 3.2)
  2. A-12.
    $\sigma$ ReLU activation (Def. 3.2)
  3. A-13.
    $L$ depth of a neural network (Def. 3.2)
  4. A-14.
    $M$ width of a neural network (Def. 3.2)
  5. A-15.
    $H_{r}$ low rank neural representation of rank $r$ (Def 3.3)
  6. A-16.
    $K$ degree of freedom of a LRNR (Def. 3.3)
  7. A-17.
    $\mathbb{W}_{r}$, $\mathbb{B}_{r}$ low dimensional weights and biases
    (3.16) $\mathbb{W}_{r}=\bigtimes_{\ell=1}^{L}(C^{\ell}\cdot\mathbb{U}^{\ell}_{r})\subset\mathbb{W}\qquad\mathbb{B}_{r}=\bigtimes_{\ell=1}^{L}(D^{\ell}\cdot\mathbb{V}^{\ell}_{r})\subset\mathbb{B}$
  8. A-18.
    LRNR (Def. 3.3)
    (3.17) $H_{r}=\left\{h(W,B)\mid W\in\mathbb{W}_{r},\ B\in\mathbb{B}_{r}\right\}$
Transported subspaces, inverse-bias trick, LRNR approximation of classical solutions (Section 4).
  1. A-19.
    $\bar{H}_{r}:=(B\cdot\Psi_{r})\circ(A\cdot\Phi_{r})^{-1}$ transported subspace (Def. 4.2)
  2. A-20.
    $\rho_{\varepsilon}$ continuous piecewise linear approximation to the threshold function (4.9)
  3. A-21.
    Inverse-bias trick for $(f,g)\in B\cdot\Psi_{r}\times A\cdot\Phi_{r}$
    (4.7) $f\circ g^{-1}(x)=c_{0}+\sum_{k}c_{k}\rho(g^{-1}(x)-x_{k})=c_{0}+\sum_{k}c_{k}\rho(x-g(x_{k}))$
  4. A-22.
    $f_{\varepsilon}\circ^{+}g_{\varepsilon}$ continuous (finitely) piecewise linear analogue of the inverse bias trick
    (4.10) $f_{\varepsilon}\circ^{+}g_{\varepsilon}(x)=\sum_{i=1}^{N}c_{i}\rho_{\varepsilon}\left(x-g_{\varepsilon}(x_{i})\right)$
  5. A-23.
    LRNR ($L=2,r=3$) approximation $h(t)\in H_{r}$ to the classical solution (Thm. 4.7)
    (4.20) $\lVert{u(t)-h(t)}\rVert_{L^{1}(\Omega_{x})}\lesssim\frac{1}{K}\left\lvert u_{0}\right\rvert_{TV(\Omega_{x})}\left(1+T\lVert{F^{\prime\prime}}\rVert_{L^{\infty}(u_{0}(\Omega_{x}))}\lVert{u_{0}^{\prime}}\rVert_{L^{\infty}(\Omega_{x})}\right)$
LRNR approximation the entropy solution (Section 5).
  1. A-24.